# Trivariate Bicycle Codes

### v1

@@ V Code layout
Physically, one can arrange all data and measurement qubits in a two-dimensional grid with periodic boundary conditions composed of two sub-lattices. For a $\mathcal{W}$ weight code, four interactions remain local, while $\mathcal{W}-4$ interactions are non-local with translational symmetry. An example is shown in Fig. 1.
Using the test in Proposition 2, we further conclude that all the weight-4 codes do not have a toric layout. After all, a weight-4 toric layout code would imply a Tanner Graph isomorphic to that of a toric code’s, and thus the TB code would be equivalent to a toric code as elaborated in Proposition 5. Thus, long-range edges are necessary to go beyond the performance of the toric code. One method to achieve long-range edges is through a twisted torus as explored in [43], or more generally a tangled toric layout which produces long-range edges, such as in Fig 9 or Ref. [44].
Definition 2 (Tangled Toric Layout).
A QLDPC code $C$ has a ($\sigma$, $\tau$)-tangled toric layout $\iff\exists\sigma\in S_{2\mu},\tau\in S_{2\lambda}$ where $S_{q}$ is the permutation group on $q$ elements, such that the Tanner Graph of $C$ has a toric layout after applying $\sigma$ along the x-direction of the periodic boundary conditions of the toric layout, similarly for $\tau$ along the y-direction.
Intuitively, a tangled toric layout means that the code behaves as a surface code in the bulk, but with long-range edges produced by modifying the PBC along the horizontal and vertical edges of the square grid. One way to formalize this is by providing bijective maps $(\sigma,\tau)$ we call tangling parameters. Such a tangled toric code would only require long-range connections along the boundary of a planar implementation, making it potentially more attractive than bi-planar architectures when taking a first step beyond the toric code. This is in stark contrast to the toric code, where we have a direct identification of opposite points of the square grid. Tangling induces long-range edges that allow one to increase the locality constant $c$ in the 2-D Bravyi-Poulin-Terhal (BPT) bound, $kd^{2}\leq cn$ [45]. For the toric code $c=1$ saturates the bound, but in our weight-4 $[[64,2,8]]$ code, choosing $c=2$ would saturate the BPT bound, verifying its enhanced performance. We provide an $O(n)$ algorithm to determine the modification to PBC of the torus. This algorithm applies more generally to two-block codes, defined as CSS QLDPC codes such that $H_{X}=[A|B]$ and $H_{Z}=[B^{T}|A^{T}]$, where A and B commute and have the same size.
@@
For any two-block CSS QLDPC code of sparsity 4, suppose it has a tangled toric layout known to have torus parameters $\mu$,$\lambda$. Then there is an $O(n)$ time algorithm to find $\sigma$, $\tau$.
The key idea of the algorithm is to embed a spanning subgraph onto the rectangular 2-D grid of size $2\mu\times 2\lambda$. The remaining edges on the boundary of the grid left unassigned to the square grid along x-direction and y-direction give us $\sigma$ and $\tau$ respectively. See details in Appendix C Proposition 9.
Figure 9: Tanner Graph of the weight-4 [[64, 2, 8]] code, as a tangled torus with layout parameters $\mu$ = l = 8, $\lambda$ = m = 4. The usual torus topology is disrupted by a tangle on the right side of the figure.
VI Discussion
In this work, we introduce TB-QLDPC codes and present compact codes with high encoding rates and strong noise suppression. Our TB-QLDPC codes are a generalisation of BB-QLDPC codes by introducing a new code variable. Moreover, by optimised search we find codes with more resource-efficient parameters than corresponding surface codes. We obtain promising codes with only weights four and five, which could offer a more hardware-friendly implementation compared to recently found weight six BB-QLDPC codes from Ref. [20, 41, 46]. For example, we find a highly compact weight-5 $[[30,4,5]]$ code which has a toric bi-planar layout and a single long-range connection, which shows promise for practical implementations on current hardware.
Further, we make progress on the properties of TB-QLDPC: We show that up to weight-6, all TB-QLPDC codes have conveniently a bi-planar structure. Further, we provide methods to check whether TB-QLDPC codes have a toric layout, which could be used to systematically search for codes with experimental-friendly layout.
To find even better low-weight codes, future research could combine our TB-QLDPC codes with subsystem codes [47, 48] and Floquet codes [49, 50, 51, 52, 53, 54]. This could develop subsystem trivariate bicycle codes and Floquet trivariate bicycle codes which offer a potential avenue for a systematic approach to reducing the weight of check operators. A potential method to construct such codes is through Floquetification via ZX calculus [55, 56].
In the future, automating the discovery of TB-QLDPC with favourable parameters, such as weight, Tanner graph thickness, encoding rate, and distance, could be achieved using machine learning techniques like reinforcement learning.
An interesting future work would involve analyzing the structure of polynomials analytically, allowing for the systematic discovery of TB codes with favorable parameters and Tanner Graph structure. Finally, it would be interesting to develop adaptive TB-QLDPC codes using quantum combs [57].
To further ease hardware implementation, future work studying larger QLDPC codes can also seek to optimise or provide heuristics for easing calculation of biplanarity. This could be done by leveraging on the symmetries in the Tanner Graph of QLDPC codes such as its toric layout or translation invariance, to obtain an algorithm that is not exponentially hard in system size. Specifically, the search space for possible subgraphs should be reduced to $\text{poly}(n)$ for an algorithm to be used in a scalable search for good bi-planar QLDPC codes.
Acknowledgements.
We thank Sergey Bravyi and Patrick Rall for a fruitful discussion in the beginning of the project. LV acknowledges the financial support by the Baden-Württemberg Foundation and the Ernst-Wilken Foundation for his research stay at CQT. This research is supported by A*STAR C230917003.
Appendix
@@ Appendix A Codes
$[[n,k,d]]$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$[[112,8,5]]$ $1/14$ 7, 8 1.8 0.0298 $2\!\times\!10^{-9}$ $z^{2}+z^{6}$ $x+x^{6}$$\times$ $\times$ $\times$ $\checkmark$
$[[64,2,8]]$ $1/32$ 8, 4 2.0 0.0767 $4\!\times\!10^{-13}$ $x+x^{2}$ $x^{3}+y$$\times$ $\times$ $\times$ $\checkmark$
$[[72,2,8]]$ $1/36$ 4, 9 1.8 0.0863 $2\!\times\!10^{-13}$ $x+y^{2}$ $x^{2}+y^{2}$$\times$ $\times$ $\times$ $\checkmark$
$[[96,2,8]]$ $1/48$ 6, 8 1.3 0.0911 $4\!\times\!10^{-16}$ $x^{5}+y^{6}$ $z+z^{4}$$\times$ $\times$ $\times$ $\checkmark$
$[[112,2,10]]$ $1/56$ 7, 8 1.8 0.097 $2\!\times\!10^{-16}$ $z^{6}+x^{5}$ $z^{2}+y^{5}$$\times$ $\times$ $\times$ $\checkmark$
$[[144,2,12]]$ $1/72$ 8, 9 2.0 0.1017 $4\!\times\!10^{-19}$ $x^{3}+y^{7}$ $x+y^{5}$$\times$ $\times$ $\times$ $\checkmark$
$[[144,2,12]]$ $1/72$ 8, 9 2.0 0.1017 $4\!\times\!10^{-19}$ $x^{3}+y^{7}$ $x^{7}+y$$\times$ $\times$ $\times$ $\checkmark$
@@
$[[n,k,d]]$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$[[30,4,5]]$ $1/7$ 3, 5 3.3 0.0437 $6\!\times\!10^{-10}$ $x+z^{4}$ $x+y^{2}+z^{2}$ $(1,2,2,3)$ $(5,3)$checkmark $\checkmark$
$[[72,4,8]]$ $1/18$ 4, 9 3.6 0.0785 $8\!\times\!10^{-14}$ $x+y^{3}$ $x^{2}+y+y^{2}$$\times$ $\times$ $\times$ $\checkmark$
$[[96,4,8]]$ $1/24$ 8, 6 2.7 0.0823 $1\!\times\!10^{-13}$ $x^{6}+x^{3}$ $z^{5}+x^{5}+y$ $(1,2,1,2)$ $(8,6)$checkmark $\checkmark$
@@
$[[n,k,d]]$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$[[30,6,4]]$ $1/5$ 5, 3 3.2 0.0234 $3\!\times\!10^{-7}$ $x^{4}+z^{3}$ $x^{4}+x+z^{4}+y$ $(1,2,1,3)$ $(5,3)$checkmark $\checkmark$
$[[48,6,6]]$ $1/8$ 4, 6 4.5 0.0495 $2\!\times\!10^{-10}$ $x^{2}+y^{4}$ $x^{3}+z^{3}+y^{2}+y$$\times$ $\times$ $\times$ $\checkmark$
$[[40,4,6]]$ $1/10$ 4, 5 3.6 0.0588 $7\!\times\!10^{-11}$ $x^{2}+y$ $y^{4}+y^{2}+x^{3}+x$$\times$ $\times$ $\times$ $\checkmark$
$[[48,4,6]]$ $1/12$ 4, 6 3.0 0.0698 $3\!\times\!10^{-11}$ $x^{3}+y^{5}$ $x+z^{5}+y^{5}+y^{2}$ $(1,2,3,4)$ $(12,2)$checkmark $\checkmark$
@@
$[[n,k,d]]$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$[[30,4,5]]$ $1/7$ 5, 3 3.3 0.0507 $5\!\times\!10^{-10}$ $x^{4}+x^{2}$ $x+x^{2}+y+z^{2}+z^{3}$ $(1,2,2,4)$ $(5,3)$checkmark $\times$
@@
Appendix B Decoding Plots
@@
We show noise suppression performance for QLDPC codes of weight six under depolarizing noise. In Figs. 1313, we show our QLDPC code in comparison with a selection of surface code of lower and higher distance than the QLDPC code to illustrate the noise performance. Since a surface code architecture would require $k$ batches of a $d\!\times\!d$ grid of qubits, the total number of physical qubits required to implement $k$ logical qubits at code distance $d$ holds for $n=kd^{2}$.
@@
Figure 10: Our weight-6 QLDPC code $[[48,6,6]]$ has a lower logical error rate than the distance-$5$ surface code while requiring three times less qubits coming with an encoding rate of $1/8$.
@@
Figure 11: Our weight-6 QLDPC code [[48, 4, 6]] has an encoding rate of 1/12. We only need a third of the number of physical qubits compared to the distance-$6$ surface code to correct up to $2$ error before a logical one would occur.
@@
Figure 12: We present a weight-6 QLDPC with an encoding rate of $10\%$ and parameters [[40, 4, 6]]. Compared to the distance-6 surface code, our code requires less physical qubits by a factor of $3.6$.
@@
Figure 13: This weight-4 QLDPC code [[112, 2, 10]] requires about half the number of physical qubits compared to the surface code.
@@
Next, we show the noise suppression performance for QLDPC codes of weight four under depolarizing noise. In Figs. 1616, we show our QLDPC code in comparison with a selection of surface code of lower and higher distance than the QLDPC code to illustrate the noise performance. Since a surface code architecture would require $k$ batches of a $d\!\times\!d$ grid of qubits, the total number of physical qubits required to implement $k$ logical qubits at code distance $d$ holds for $n=kd^{2}$.
@@
Figure 14: Our weight-4 QLDPC code [[72, 2, 8]] offers a high code distance while just using almost half of the number of physical qubits than the surface code would require.
@@
Figure 15: Our weight-4 QLDPC code [[96, 2, 8]] provides good noise suppression characteristics and requires less physical qubits than the equivalent surface by $33\%$.
@@
Figure 16: Our weight-4 QLDPC code $[[64,2,8]]$ encodes $2$ logical qubits at only half the resources needed for a surface code implementation.
@@
Finally, we show the suppression performance for QLDPC codes of weight five under depolarizing noise. In Figs. 1818, we show our QLDPC code in comparison with a selection of surface code of lower and higher distance than the QLDPC code to illustrate the noise performance. Since a surface code architecture would require $k$ batches of a $d\!\times\!d$ grid of qubits, the total number of physical qubits required to implement $k$ logical qubits at code distance $d$ holds for $n=kd^{2}$.
@@
Figure 17: Our weight-5 QLDPC code $[[72,4,8]]$ shows better noise performance than the distance-$7$ surface code and is more qubit efficient by a factor of $2.7$. We find a fitted code distance of $8.3$.
@@
Figure 18: With our weight-5 QLDPC code $[[96,4,8]]$ we provide a mix out of high code distance and moderate encoding rate. Further, an implementation requires less physical qubits than the distance-8 surface code by almost a factor of $2.7$.
Appendix C Tanner Graph Details

### v2

@@ Abstract
For a weight-$\mathcal{W}$ TB code, four interactions remain local, while $\mathcal{W}-4$ interactions are non-local with translational symmetry. An example is shown in Fig. 1. The key property allowing this is the cyclic property of $H_{X}$ and $H_{Z}$ for such codes, elaborated in the Supplemental Material (SM) [25]. An analogous proposition holds for MB codes with a toric layout as well, by repeating the arguments in higher spatial dimensions.
Using the test in Proposition 2, we further conclude that all the weight-4 codes do not have a toric layout. However, they can possess a tangled toric layout (see Def 9), such as in Fig 4 by using the tool from Ref. [26].
Intuitively, a tangled toric layout means that the code behaves as a surface code in the bulk, but with long-range edges produced by modifying the PBC along the horizontal and vertical edges of the square grid. This minor modification makes it potentially more attractive than bi-planar architectures when taking a first step beyond the toric code. In the Bravyi-Poulin-Terhal (BPT) bound $kd^{2}\leq cn$ [27], for the toric code the locality constant $c=1$ saturates the bound, but in our weight-4 $\llbracket 64,2,8\rrbracket$ code we would need $c=2$, verifying its enhanced performance. We provide an $O(n)$ algorithm to determine the tangling parameters ($\sigma$, $\tau$). This algorithm applies more generally to two-block codes, defined as CSS QLDPC codes such that $H_{X}=[A|B]$ and $H_{Z}=[B^{T}|A^{T}]$, where A and B commute and have the same size.
@@
For any two-block CSS QLDPC code of sparsity 4 with equal X and Z degree per data qubit (2 each), suppose it has a tangled toric layout known to have torus parameters $\mu$,$\lambda$. Then there is an $O(n)$ time algorithm to find $\sigma$, $\tau$.
The key idea of the algorithm is to embed a spanning subgraph onto the rectangular 2-D grid of size $2\mu\times 2\lambda$. The unassigned edges on the grid give us $\sigma$ and $\tau$ respectively. Details in Proposition 10.
Figure 4: Tanner Graph of the weight-4 $\llbracket 64,2,8\rrbracket$ code, as a tangled torus with layout parameters $\mu=l=8,\lambda=m=4$. The usual torus topology is disrupted by a tangle on the right side of the figure.
Discussion.—  In this work, we introduce MB-QLDPC codes as multi-variable generalisation of BB-QLDPC codes, which have only two variables. We study TB-QLDPC with three variables in detail and present compact codes with high encoding rates and strong noise suppression. Moreover, by optimised search we find codes with more resource-efficient parameters than corresponding surface codes. We obtain promising codes with only weights four and five, which could offer a more hardware-friendly implementation compared to recently found weight six BB-QLDPC codes from Ref. [20, 24, 28]. For example, we find a highly compact weight-5 $\llbracket 30,4,5\rrbracket$ code which has a toric bi-planar layout and a single long-range connection, which shows promise for practical implementations on current hardware. Further, we make progress on the properties of TB-QLDPC codes: We show that up to weight-6, all TB-QLPDC codes have conveniently a bi-planar structure. Additionally, we provide methods to check whether TB-QLDPC codes have a toric layout, which could be used to systematically search for codes with experimental-friendly layout, significantly reducing the search space of codes. Future research directions are discussed in SM E.
Acknowledgments.—  We thank Sergey Bravyi and Patrick Rall for a fruitful discussion in the beginning of the project. LV acknowledges the financial support by the Baden-Württemberg Foundation and the Ernst-Wilken Foundation for his research stay at CQT. This research is supported by A*STAR C230917003.
Supplemental Material
@@ Appendix A Codes overview
$\llbracket n,k,d\rrbracket$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$\llbracket 112,8,5\rrbracket$ $1/14$ 7, 8 1.8 0.0298 $2\!\times\!10^{-9}$ $z^{2}+z^{6}$ $x+x^{6}$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 64,2,8\rrbracket$ $1/32$ 8, 4 2.0 0.0767 $4\!\times\!10^{-13}$ $x+x^{2}$ $x^{3}+y$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 72,2,8\rrbracket$ $1/36$ 4, 9 1.8 0.0863 $2\!\times\!10^{-13}$ $x+y^{2}$ $x^{2}+y^{2}$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 96,2,8\rrbracket$ $1/48$ 6, 8 1.3 0.0911 $4\!\times\!10^{-16}$ $x^{5}+y^{6}$ $z+z^{4}$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 112,2,10\rrbracket$ $1/56$ 7, 8 1.8 0.097 $2\!\times\!10^{-16}$ $z^{6}+x^{5}$ $z^{2}+y^{5}$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 144,2,12\rrbracket$ $1/72$ 8, 9 2.0 0.1017 $4\!\times\!10^{-19}$ $x^{3}+y^{7}$ $x+y^{5}$$\times$ $\times$ $\times$ $\checkmark$
@@
$\llbracket n,k,d\rrbracket$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$\llbracket 30,4,5\rrbracket$ $1/7$ 3, 5 3.3 0.0437 $6\!\times\!10^{-10}$ $x+z^{4}$ $x+y^{2}+z^{2}$ $(1,2,2,3)$ $(5,3)$checkmark $\checkmark$
$\llbracket 72,4,8\rrbracket$ $1/18$ 4, 9 3.6 0.0785 $8\!\times\!10^{-14}$ $x+y^{3}$ $x^{2}+y+y^{2}$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 96,4,8\rrbracket$ $1/24$ 8, 6 2.7 0.0823 $1\!\times\!10^{-13}$ $x^{6}+x^{3}$ $z^{5}+x^{5}+y$ $(1,2,1,2)$ $(8,6)$checkmark $\checkmark$
@@
$\llbracket n,k,d\rrbracket$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$\llbracket 30,6,4\rrbracket$ $1/5$ 5, 3 3.2 0.0234 $3\!\times\!10^{-7}$ $x^{4}+z^{3}$ $x^{4}+x+z^{4}+y$ $(1,2,1,3)$ $(5,3)$checkmark $\checkmark$
$\llbracket 48,6,6\rrbracket$ $1/8$ 4, 6 4.5 0.0495 $2\!\times\!10^{-10}$ $x^{2}+y^{4}$ $x^{3}+z^{3}+y^{2}+y$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 40,4,6\rrbracket$ $1/10$ 4, 5 3.6 0.0588 $7\!\times\!10^{-11}$ $x^{2}+y$ $y^{4}+y^{2}+x^{3}+x$$\times$ $\times$ $\times$ $\checkmark$
$\llbracket 48,4,6\rrbracket$ $1/12$ 4, 6 3.0 0.0698 $3\!\times\!10^{-11}$ $x^{3}+y^{5}$ $x+z^{5}+y^{5}+y^{2}$ $(1,2,3,4)$ $(12,2)$checkmark $\checkmark$
@@
$\llbracket n,k,d\rrbracket$ $r$ $l,m$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{L}(10^{-4})$ $A$ $B$ $i,j,g,h$ $\mu,\lambda$ toric bi-planar
$\llbracket 30,4,5\rrbracket$ $1/7$ 5, 3 3.3 0.0507 $5\!\times\!10^{-10}$ $x^{4}+x^{2}$ $x+x^{2}+y+z^{2}+z^{3}$ $(1,2,2,4)$ $(5,3)$checkmark $\times$
@@
Appendix B Selected Codes of different weight
Weight-4.—  In Fig. 6, we show the physical error rate $p$ against the logical error rate $p_{\text{L}}$ for our weight-4 QLDPC code $\llbracket 112,8,5\rrbracket$. As comparison, we also show surface codes with the same number of logical qubits and similar distance $d$. We find that our code has comparable error suppression than the $\llbracket 200,8,5\rrbracket$ surface code, while requiring only about half the number of physical qubits. We also fit the logical error rates with Eq. (4), where for our code with find a fitted code distance $d_{\text{fit}}=5.9$.
Weight-5.—  Next, in Fig. 6 we study the error suppression of our $\llbracket 30,4,5\rrbracket$ code, which is a weight-$5$ TB code, who’s layout is shown in Fig. 1. Our code suppresses logical errors at a physical error rate $p=10^{-3}$ by more than three orders of magnitude, allows for a toric bi-planar layout and has about three times higher encoding rate compared to the $\llbracket 100,4,5\rrbracket$ surface code which has similar error suppression.
Figure 5: Logical error rate $p_{\text{L}}$ against physical error rate $p$ for our weight-4 code $\llbracket 112,8,5\rrbracket$, and surface codes with similar code parameters. The error curves are fitted with Eq. (4), yielding a fitted code distance of $d_{\text{fit}}=5.9$.
Figure 6: Logical error rate $p_{\text{L}}$ against physical error rate $p$ for our weight-5 code $\llbracket 30,4,5\rrbracket$. We show surface codes with comparable code parameters as reference. We find $d_{\text{fit}}=6.1$ for our QLDPC code.
Figure 7: Logical error rate of our weight-6 code $\llbracket 30,6,4\rrbracket$ against comparable surface codes. We find $d_{\text{fit}}=4.1$.
Figure 8: Our weight-7 code $\llbracket 30,4,5\rrbracket$ compared against surface codes with the same number of logical qubits. We find $d_{\text{fit}}=6.1$.
Weight-6.—  In Fig. 8, we present our weight-6 $\llbracket 30,6,4\rrbracket$ code which has a high encoding rate of $1/5$ and find it to have a toric layout with bi-planar structure. We find comparable error suppression scaling as the distance $3$ surface code. At physical error rate $p=10^{-3}$, we find a logical error rate of $3.5\times 10^{-5}$ corresponding to a logical noise suppression by two orders of magnitude. We note that a surface code with distance-$4$ and the same number of logical qubits requires $96$ physical qubits, making our code three times more qubit efficient.
Weight 7.—  Now, in Fig. 8, we study our weight-7 code with parameters $\llbracket 30,4,5\rrbracket$. Note that we also found a weight-$5$ code with the the same code parameters. We find that our weight-7 $\llbracket 30,4,5\rrbracket$ code has similar error suppression than the $\llbracket 100,4,5\rrbracket$ surface code.
Appendix C Decoding performance
@@
We show noise suppression performance for our QLDPC codes of weight six under depolarizing noise. In Figs. 1212, we show our codes in comparison with a selection of surface code of lower and higher distance than the QLDPC code to illustrate the noise performance. Since a surface code architecture would require $k$ batches of a $d\!\times\!d$ grid of qubits, the total number of physical qubits required to implement $k$ logical qubits at code distance $d$ reads $n=kd^{2}$.
@@
Figure 9: Our weight-6 code $\llbracket 48,6,6\rrbracket$ has a lower logical error rate than the distance-$5$ surface code while requiring three times less qubits coming with an encoding rate of $1/8$.
@@
Figure 10: Our weight-6 code $\llbracket 48,4,6\rrbracket$ has an encoding rate of 1/12. We only need a third of the number of physical qubits compared to the distance-$6$ surface code to correct up to $2$ error before a logical one would occur.
@@
Figure 11: We present a weight-6 with an encoding rate of $10\%$ and parameters $\llbracket 40,4,6\rrbracket$. Compared to the distance-6 surface code, our code requires less physical qubits by a factor of $3.6$.
@@
Figure 12: This weight-4 code $\llbracket 112,2,10\rrbracket$ requires about half the number of physical qubits compared to the surface code.
@@
Next, we show the noise suppression performance for QLDPC codes of weight four under depolarizing noise. In Figs. 1515, we show our codes in comparison with a selection of surface code of lower and higher distance than the QLDPC code to illustrate the noise performance. Since a surface code architecture would require $k$ batches of a $d\!\times\!d$ grid of qubits, the total number of physical qubits required to implement $k$ logical qubits at code distance $d$ reads $n=kd^{2}$.
@@
Figure 13: Our weight-4 code $\llbracket 72,2,8\rrbracket$ offers a high code distance while just using almost half of the number of physical qubits than the surface code would require.
@@
Figure 14: Our weight-4 code $\llbracket 96,2,8\rrbracket$ provides good noise suppression characteristics and requires less physical qubits than the equivalent surface by $33\%$.
@@
Figure 15: Our weight-4 code $\llbracket 64,2,8\rrbracket$ encodes two logical qubits at only half the resources needed for a surface code implementation.
@@
Finally, we show the suppression performance for QLDPC codes of weight five under depolarizing noise. In Figs. 1717, we show our code in comparison with a selection of surface code of lower and higher distance than the code to illustrate the noise performance. Since a surface code architecture would require $k$ batches of a $d\!\times\!d$ grid of qubits, the total number of physical qubits required to implement $k$ logical qubits at code distance $d$ reads $n=kd^{2}$.
@@
Figure 16: Our weight-5 code $\llbracket 72,4,8\rrbracket$ shows better noise performance than the distance-$7$ surface code and is more qubit efficient by a factor of $2.7$. We find a fitted code distance of $8.3$.
@@
Figure 17: With our weight-5 code $\llbracket 96,4,8\rrbracket$ we provide a mix out of high code distance and moderate encoding rate. Further, an implementation requires less physical qubits than the distance-8 surface code by almost a factor of $2.7$.
Appendix D Background
QEC codes are essential for executing high-fidelity computations on noisy quantum computers. Without QEC, errors accumulate during circuit execution, yielding an unreliable output. Stabilizer codes [29, 30] represent a class of QEC codes defined by their stabilizer group, which is an abelian subgroup of the Pauli group on $n$ qubits that keeps the codespace invariant. The codespace of a stabilizer code is equivalently described as the joint $+1$-eigenspace of the stabilizer generators $S=\langle S_{1},S_{2},...,S_{r}\rangle$. For a $\llbracket n,k,d\rrbracket$ code with $n$ physical qubits, $k$ logical qubits, and distance $d$, the number of linearly independent generators is $r=n\!-\!k$. A stabilizer code qualifies as a QLDPC code if each qubit supports a constant number of stabilizer generators and each generator has a constant weight independent of $n$. A stabilizer code is classified as a CSS code [31, 7] if each generator is a tensor product of $X$ and $I$ or a tensor product of $Z$ and $I$. Although the surface code is a QLDPC code, its encoding rate $k/n$ approaches zero as $n$ approaches infinity, contributing to its high overhead. In contrast, alternative QLDPC codes maintain asymptotically constant encoding rates while preserving or improving the $d=\mathcal{O}(\sqrt{n})$ distance scaling of the surface code [32, 33, 16, 34, 35, 36, 37].
Appendix E Outlook
To find even better low-weight codes, future research could combine our MB-QLDPC codes with subsystem codes [46, 47] and Floquet codes [48, 49, 50, 51, 52, 53]. This could develop subsystem trivariate bicycle codes and Floquet trivariate bicycle codes which offer a potential avenue for a systematic approach to reducing the weight of check operators. A potential method to construct such codes is through Floquetification via ZX calculus [54, 55]. In the future, automating the discovery of MB-QLDPC with favourable parameters, such as weight, Tanner graph thickness, encoding rate, and distance, could be achieved using machine learning techniques like reinforcement learning. An interesting future work would involve analyzing the structure of polynomials analytically, allowing for the systematic discovery of MB codes with favorable parameters and Tanner Graph structure. Finally, it would be interesting to develop adaptive MB-QLDPC codes using quantum combs [56]. To further ease hardware implementation, future work studying larger QLDPC codes can also seek to optimise or provide heuristics for easing calculation of biplanarity. This could be done by leveraging on the symmetries in the Tanner Graph of QLDPC codes such as its toric layout or translation invariance, to obtain an algorithm that is not exponentially hard in system size. Specifically, the search space for possible subgraphs should be reduced to $\text{poly}(n)$ for an algorithm to be used in a scalable search for good bi-planar QLDPC codes.
Appendix F Tanner Graph Details